OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence?

OpenStudy (anonymous):

first term n = 1 second n=2 third n=3 fourth n =4

OpenStudy (anonymous):

1. 4*1 -1 =? 2. 4*2 -1 =? 3. 4*3 -1 =? 4. 4*4 -1 =?

OpenStudy (anonymous):

wait..im sorry they gave me answer choices and those are none of em

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

no i still dont understand

OpenStudy (anonymous):

first term impliest that it's the first number in sequence, therefore you use 1 instead of n

OpenStudy (anonymous):

second term means second number in sequence, therefore you use 2 instead of n

OpenStudy (anonymous):

ohhh i get it

OpenStudy (anonymous):

does answers match?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the second answer choice

OpenStudy (anonymous):

okay, goodnight then :)

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can you help me on a different question

OpenStudy (anonymous):

its way different